OpenStudy (amenah8):

Shakira plans to build a 4 side-by-side rectangular enclosure whenever/wherever she wants around her garden. SHE DECIDES TO USE HER HOUSE AS ONE SIDE OF THE GARDEN. She has 1800 ft of fence. What would be the length and width of the garden for maximum AREA? What is the max area?

OpenStudy (anonymous):

calc or no? doesn't make a difference, just asking

OpenStudy (amenah8):

yep. calculus :)

OpenStudy (danjs):

draw it, and total the perimeter using a variable for length and width

OpenStudy (anonymous):

not needed but we can do it that way

OpenStudy (anonymous):

first off, although i am not supposed to give a direct answer, lets answer it first

OpenStudy (anonymous):

use half of the fence for the side opposite the house, split the remaining half for the other two sides there, i said it, now lets find it

OpenStudy (anonymous):



OpenStudy (amenah8):

right

OpenStudy (anonymous):

how much is left for the part of the fence not labelled?

OpenStudy (amenah8):

A=lw so A=(1800-2x)(x) ??

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

now for the lets not use calculus part, because it is unnecessary \[A(x)=1800x-2x^2\] is a parabola that opens down the max is at the vertex use \(-\frac{b}{2a}\) to find it

OpenStudy (danjs):

i thought the thing was 4 rectangles side by side p=5x+y,, maybe not

OpenStudy (anonymous):

or if you are a slave to calc, take the derivative, set it equal to zero and solve you will still get \(-\frac{b}{2a}\)

OpenStudy (anonymous):

oh crap i should learn how to read !!!

OpenStudy (anonymous):

ignore all i said, @DanJS pointed out the correct question

OpenStudy (danjs):

not sure, you may be right, enclosure is singular, not 4 of em..

OpenStudy (amenah8):

x-450

OpenStudy (amenah8):

sorry, x=450

OpenStudy (danjs):

either way, same process to solve , sorry

OpenStudy (anonymous):

i am not sure what a "4 side by side" rectangular enclosure is maybe it is this 

OpenStudy (anonymous):

no i am sorry, i didn't read carefully enough

OpenStudy (amenah8):

right. that's what i was thinking

OpenStudy (anonymous):

then it is \(A=x(1800-5x)\) and the solution is different

OpenStudy (amenah8):

so x is not 450?

OpenStudy (anonymous):

not any more

OpenStudy (anonymous):

now \[A=1800x-5x^2\] so it is different

OpenStudy (anonymous):

i mean the work is the same, but the answer is different

OpenStudy (amenah8):

ok, one moment while i try to solve for x

OpenStudy (anonymous):

ok, take the derivative first of course, then set that equal to zero

OpenStudy (amenah8):

take the derivative of 1800x-5x^2?

OpenStudy (anonymous):

yes

OpenStudy (amenah8):

A(x)= 1800x - 5x^2 --> 1st D= 1800-10x

OpenStudy (amenah8):

x= 180

OpenStudy (anonymous):

yes

OpenStudy (amenah8):

so now, A=(180)(w)

OpenStudy (anonymous):

\[ A=180(1800-5\times 180)\]

OpenStudy (amenah8):

ah, so plug in 180 for x into x(1800-5x)

OpenStudy (anonymous):

yes

OpenStudy (amenah8):

-1296000?

OpenStudy (anonymous):

no

OpenStudy (amenah8):

oh, i see what i did wrong. it's 162000

OpenStudy (anonymous):

looks good

OpenStudy (amenah8):

dimensions are 180 x 900

OpenStudy (anonymous):

yes i think so

OpenStudy (amenah8):

THANK you so much! this really helped me. :)

OpenStudy (anonymous):

yw